A flag pole 20 feet high casts a shadow of 5 feet. At the same time, a nearby building casts a 20-foot shadow. How many feet high is the building? ( the ground is level and both the building and the flagpole are perpendicular to the ground)
Answer by mmm4444bot(95)

Hello There:
This problem can be solved by using similar triangles. Each triangle is a right triangle.
The shadow of the flag pole is the base of the triangle, and its height is the other leg.
The shadow of the building is the base of the other triangle, and its height is unknown. Let's call the height "x".
Now we can write the equation relating the ratios of height to base to each other.
20/5 = x/20
Cross multiply to get rid of the fractions.
400 = 5*x
x = 80
The building is 80 feet tall.
~ Mark